Three charges -q,+q and -q are placed at the corners of an equilateral triangle of side `a` . The resultant electric force on a charge +q placed at the centroid O of the triangle is

Grade:10

2 Answers

Suraj Prasad IIT Patna
7 years ago
See for these type of questions. Draw the triangle and put charges at the three vertex and try to get the resultant force at the centroid by using the formula:

F = K q1 q2/ r^2

You need to take care of the direction of the force as well.
S Reddy priya
6 years ago
Distance between -q to charge at centroid q is a/sqrt3. Force b/w them is F1=k×(-3q^2/a^2). Similarly F2=k×(-3q^2/a^2),F3=k×(3q^2/a^2). Resultant force F=F1+F2+F3 =K×(3q^2/a^2)